[Brahms-l] Shift report 20050316 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Mar 16 2005 - 16:03:38 EST
Shifter leader was rd
Shifters were 


1. There was a controlled acces for ~3 hours followed by a power dip. We only had 2 hours of beam.

2. I collected 30K trigger 2 at 8 degrees 1219A

3. Tried to fix the C4 channels that cannot reach nominal HV but failed, they continue to run ~100V below demand.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Mar 16 16:04:14 2005

This archive was generated by hypermail 2.1.8 : Wed Mar 16 2005 - 16:04:29 EST